OpenStudy (anonymous):

The bones of a Awesomasaurua were calculated to have 2% of their carbon-14 remaining? If the half life of carbon-14 is 5760 years, how long ago did it die? --------------------------------------------------------------------------------

OpenStudy (anonymous):

\[(\frac{1}{2})^{\frac{t}{5760}}=.02\] solve for t

OpenStudy (anonymous):

\[\frac{t}{5760}=\frac{\ln(.02)}{\ln(.5)}\] \[t=5760\times \frac{\ln(.02)}{\ln(.5)}\]

OpenStudy (anonymous):

ok,ty

OpenStudy (anonymous):

yw